OpenStudy (crashonce):

For an AP, the sum of the first 100 and the sum of the first 10 terms are 10 and 100 respectively. What is the sum of the first 110 terms of the AP.

OpenStudy (crashonce):

@wio a easy one

OpenStudy (anonymous):

Hmmm, yeah, it's not too hard. Do you need help?

OpenStudy (crashonce):

ye just show me the method lol

OpenStudy (crashonce):

i keep getting the wrong answer

OpenStudy (anonymous):

First, do you know the formula for an AP?

OpenStudy (anonymous):

\[ a_n = a_1+(n-1)d \]

OpenStudy (anonymous):

\[ S_n=\frac{n}{2}\left(\frac{a_1+a_n}{2}\right) \]

OpenStudy (crashonce):

ye i know

OpenStudy (anonymous):

Combine to get: \[ \frac{n}{2}\left(\frac{2a_1+(n-1)d}{2}\right) \]

OpenStudy (anonymous):

Plugging in \(n\) and \(S_n\) will give you two formulas, which you can use to solve for \(a_1\) and \(d\).

OpenStudy (crashonce):

what is your answer

OpenStudy (anonymous):

No, I refuse to give you the answer. I have given you the means to solve it.

OpenStudy (crashonce):

i cant get the right answer, mine has heaps of fractions

OpenStudy (anonymous):

When \(n=100\), \(S_n=10\)

OpenStudy (anonymous):

So\[ \frac{100}{2}\left(\frac{2a_1+(100-1)d}{2}\right) = 10 \]

OpenStudy (anonymous):

When \(n=10\), \(S_n=100\).

OpenStudy (anonymous):

So\[ \frac{10}{2}\left(\frac{2a_1+(10-1)d}{2}\right) = 100 \]

OpenStudy (crashonce):

is d -11/25

OpenStudy (crashonce):

my answer is -2310

OpenStudy (crashonce):

the options are 90, -90, 110, -110, 100

OpenStudy (crashonce):

@wio help

OpenStudy (zarkon):

the first thing you need to do is to fix wio's formula. it is a little off

OpenStudy (anonymous):

Whoops, I divided by two twice.

OpenStudy (crashonce):

yea thats why

OpenStudy (freethinker):

"what is your answer?" laughing out loud

OpenStudy (crashonce):

k if u could help me, that would be greatly appreciated

ganeshie8 (ganeshie8):

\[S_n = \frac{n}{2}(2a_1 + (n-1)d)\] \[S_{100} = 10 \implies 50(2a_1 + 99d)=10\tag{1}\] \[S_{10} = 100\implies 5(2a_1+9d) = 100\tag{2}\] two equations and two unknowns, you can solve them

OpenStudy (crashonce):

i got 0/..........

OpenStudy (crashonce):

@ganeshie8 i still can't get it please help me

OpenStudy (crashonce):

can u check that d = -1/5 and a1 = 10.9

ganeshie8 (ganeshie8):

the idea is to fist solve \(a_1\) and \(d\), then use thes to find S_100

OpenStudy (crashonce):

but mine is still wrong

ganeshie8 (ganeshie8):

Ohk you have already solved a and d

OpenStudy (crashonce):

but my answer is 0. it is multiple choice and the options are 90, -90, 110, -110, -100

ganeshie8 (ganeshie8):

yeah it looks messy, wish there is a better method

OpenStudy (crashonce):

did u get 0 as well?

ganeshie8 (ganeshie8):

a1 = 10.99 d = -0.22

ganeshie8 (ganeshie8):

use them

OpenStudy (crashonce):

-110...

OpenStudy (crashonce):

i must be really bad but thanks both of u